OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation quantity x plus 1 squared divided by 225 minus the quantity of y plus 5 squared divided by 400 = 1. I know what a,b,c are I just don't know how to apply it

OpenStudy (anonymous):

@satellite73 @AccessDenied @sammixboo # @dan815

OpenStudy (anonymous):

- (y+5)^2 divided by 400 =1

OpenStudy (anonymous):

no actually \[\frac{(x+1)^2}{22 5}-\frac{(y+5)^2}{400}=1\]

OpenStudy (anonymous):

Yes thats it

OpenStudy (anonymous):

whew you got the center right?

OpenStudy (anonymous):

(-1,-5)?

OpenStudy (anonymous):

and also \(a==15,b=20\)

OpenStudy (anonymous):

c=25

OpenStudy (anonymous):

yeah that is right do you know what it looks like?q

OpenStudy (anonymous):

no ;

OpenStudy (anonymous):



OpenStudy (anonymous):

you need that to figure out all the rest

OpenStudy (anonymous):

horizontal

OpenStudy (anonymous):

nope first one because the \(x^2\) term comes first

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so you need the vertices right? since we now know it looks like the first one, they are 15 units to the left and right of \((-1,-5)\)

OpenStudy (anonymous):

-16,-20?

OpenStudy (anonymous):